A motor cycle is accelerated from rest to 60m/s in 16 seconds. If the engine force required to achieve this 1200N and effects due to friction are ignored, calculate the mass of the motor cycle.


1
Expert's answer
2022-03-07T10:18:53-0500

The acceleration of the motor cycle

"a=\frac{\Delta v}{\Delta t}=\frac{60}{16}=3.75\:\rm m/s^2"

The Newton's second law says

"ma=F"

Hence, the mass of motor cycle

"m=\frac{F}{a}=\frac{1200}{3.75}=320\:\rm kg"